OpenStudy (anonymous):

How long would it take $2,500 to grow to $9,700 if the annual rate is 4.1% and interest in compounded monthly? If $8,000 is placed in an account with an annual interest rate of 4%, how long will it take the amount to quadruple if the interest is compounded annually? Round your answer to two decimal places. The bacteria in a certain culture double every 6.5 hours. The culture has 7,500 bacteria at the start. How many bacteria will the culture contain after 4 hours?

jimthompson5910 (jim_thompson5910):

How long would it take $2,500 to grow to $9,700 if the annual rate is 4.1% and interest in compounded monthly? A = P(1+r/n)^(n*t) 9700 = 2500(1+0.041/12)^(12*t) 9700 = 2500(1+0.003416667)^(12*t) 9700 = 2500(1.003416667)^(12*t) 9700/2500 = (1.003416667)^(12*t) 3.88 = (1.003416667)^(12*t) (1.003416667)^(12*t) = 3.88 log( (1.003416667)^(12*t) ) = log( 3.88 ) 12*t*log( 1.003416667 ) = log( 3.88 ) I'll let you finish

OpenStudy (anonymous):

Got it thank you. I got the second one as well. I just need help with the last one.

jimthompson5910 (jim_thompson5910):

A = P*2^(6.5/t) A = 7500*2^(6.5/t) A = 7500*2^(6.5/4) A = 7500*2^(1.625) .. .. A = ??

OpenStudy (anonymous):

23,133

OpenStudy (anonymous):

Thank you! :)

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (anonymous):

Thanks so much for your help!

jimthompson5910 (jim_thompson5910):

yw